# KARMA: Leveraging Multi-Agent LLMs for Automated Knowledge Graph Enrichment

### v1

@@ 6 Conclusion
Limitations
@@
Ethical Impact
KARMA holds significant potential for automating the enrichment of knowledge graphs, particularly in complex fields like healthcare and biomedical research. However, as with any automated system, there are ethical concerns, particularly regarding bias in LLMs. Since LLMs are trained on vast and diverse datasets, they may inadvertently reflect outdated or biased information, leading to incorrect associations in the knowledge graph. Although KARMA incorporates mechanisms for verification and conflict resolution, human oversight remains essential to ensure the accuracy of critical knowledge. Additionally, considerations around data privacy are important, especially when dealing with sensitive research data. Going forward, balancing automation with human judgment will be crucial to ensuring the system operates responsibly and adheres to ethical standards. With careful attention to these challenges, KARMA has the potential to be a transformative tool for advancing knowledge while minimizing unintended consequences.
Appendix
Appendix A Detailed propmts for KARMA agents
@@
A.1 Function summaries of different agents
@@
  • Central Controller Agent (CCA): Orchestrates task scheduling and resource allocation. Uses LLM-based utility scoring and multi-armed bandit-inspired exploration to prioritize tasks (e.g., ingestion vs. conflict resolution) while balancing agent workloads.
@@
A.2 Additional Notes on Prompt Engineering
@@
A.3 Ingestion Agent (IA) Prompt
@@
A.4 Reader Agent (RA) Prompt
@@
A.5 Summarizer Agent (SA) Prompt
@@
A.6 Entity Extraction Agent (EEA) Prompt
@@
A.7 Relationship Extraction Agent (REA) Prompt
@@
A.8 Schema Alignment Agent (SAA) Prompt
@@
A.9 Conflict Resolution Agent (CRA) Prompt
@@
A.10 Evaluator Agent (EA) Prompt for Confidence
@@
A.11 Evaluator Agent (EA) Prompt for Clarity
@@
A.12 Evaluator Agent (EA) Prompt for Relevance
@@
Appendix B Examples of extracted knowledge graphs
B.1 Knowledge graph from Genomics articles (Generate using GPT-4o, Examples)
Genomics Knowledge Graph Triples Key: Conf = Confidence, Rel = Relevance, Clr = Clarity EGFR $\xrightarrow{\text{causes}}$ lung adenocarcinoma (Conf: 0.75, Rel: 0.40, Clr: 0.60) EGFR $\xrightarrow{\text{causes}}$ non - small cell lung cancer (Conf: 0.85, Rel: 0.30, Clr: 0.70) T790M $\xrightarrow{\text{causes}}$ lung adenocarcinoma (Conf: 0.98, Rel: 0.20, Clr: 0.70) T790M $\xrightarrow{\text{causes}}$ non - small cell lung cancer (Conf: 0.95, Rel: 0.20, Clr: 0.80) EGFR $\xrightarrow{\text{activates}}$ EG (Conf: 0.75, Rel: 0.20, Clr: 0.99) EGFR $\xrightarrow{\text{causes}}$ proliferation (Conf: 0.85, Rel: 0.50, Clr: 0.60) MTX - 531 $\xrightarrow{\text{treats}}$ HNSCC (Conf: 0.99, Rel: 0.80, Clr: 0.50) MTX - 531 $\xrightarrow{\text{used\_in}}$ PDX (Conf: 0.75, Rel: 0.80, Clr: 0.50) PIK3CA mutations $\xrightarrow{\text{associated\_with}}$ HNSCC (Conf: 0.85, Rel: 0.70, Clr: 0.99) MTX - 531 $\xrightarrow{\text{treats}}$ PIK3CA mutations (Conf: 0.65, Rel: 0.40, Clr: 0.99) MCLA - 158 $\xrightarrow{\text{targets}}$ EGFR (Conf: 0.75, Rel: 0.30, Clr: 0.99) EGFR $\xrightarrow{\text{interacts\_with}}$ MCLA - 158 (Conf: 0.75, Rel: 0.20, Clr: 0.80) PDOs $\xrightarrow{\text{used\_in}}$ biobank (Conf: 0.75, Rel: 0.50, Clr: 0.50) Retinoic acid $\xrightarrow{\text{disrupts}}$ autocrine growth pathway (Conf: 0.75, Rel: 0.80, Clr: 0.60) Retinoic acid $\xrightarrow{\text{interacts\_with}}$ nuclear retinoic acid receptors (Conf: 0.95, Rel: 0.80, Clr: 0.99) nuclear retinoic acid receptors $\xrightarrow{\text{regulate}}$ gene transcription (Conf: 0.95, Rel: 0.70, Clr: 0.50) EGFR $\xrightarrow{\text{interacts\_with}}$ Cys797 (Conf: 0.95, Rel: 0.80, Clr: 0.99)
@@ B.1 Knowledge graph from Genomics articles (Generate using GPT-4o, Examples)
B.2 Knowledge graph from Proteomics articles (Generate using DeepSeek-V3, Examples)
Proteomics Knowledge Graph Triples Key: Conf = Confidence, Rel = Relevance, Clr = Clarity p53 $\xrightarrow{\text{induces}}$ cell-cycle arrest (Conf: 0.95, Rel: 0.70, Clr: 0.85) p53 $\xrightarrow{\text{induces}}$ apoptosis (Conf: 0.95, Rel: 0.70, Clr: 0.85) mutant p53 $\xrightarrow{\text{causes}}$ chemotherapy resistance (Conf: 0.85, Rel: 0.85, Clr: 0.85) MDM2 $\xrightarrow{\text{interacts\_with}}$ p53 (Conf: 0.95, Rel: 0.20, Clr: 0.90) PRIMA-1 $\xrightarrow{\text{induces}}$ apoptosis (Conf: 0.85, Rel: 0.70, Clr: 0.85) NOS2 $\xrightarrow{\text{associated\_with}}$ cancers (Conf: 0.85, Rel: 0.70, Clr: 0.60) NOS2 inhibitors $\xrightarrow{\text{inhibits}}$ NOS2 (Conf: 0.95, Rel: 0.80, Clr: 0.90) NOS2 inhibitors $\xrightarrow{\text{reduces}}$ tumor growth (Conf: 0.85, Rel: 0.75, Clr: 0.85) NO $\xrightarrow{\text{induces}}$ VEGF (Conf: 0.78, Rel: 0.70, Clr: 0.70) NO $\xrightarrow{\text{induces}}$ neovascularization (Conf: 0.45, Rel: 0.70, Clr: 0.75) NOS2 inhibitors $\xrightarrow{\text{has\_therapeutic\_potential}}$ p53-mutant cancers (Conf: 0.78, Rel: 0.75, Clr: 0.85) tumor progression $\xrightarrow{\text{dependent\_on}}$ p53 (Conf: 0.85, Rel: 0.70, Clr: 0.85) NOS2 $\xrightarrow{\text{promotes}}$ tumor growth (Conf: 0.85, Rel: 0.85, Clr: 0.75) NOS2 $\xrightarrow{\text{produces}}$ nitric oxide (NO) (Conf: 0.95, Rel: 0.90, Clr: 0.90) hypoxia $\xrightarrow{\text{regulates}}$ iNOS expression (Conf: 0.85, Rel: 0.70, Clr: 0.85) iNOS expression $\xrightarrow{\text{influences}}$ endothelial integrity (Conf: 0.85, Rel: 0.70, Clr: 0.75) sulindac sulfide $\xrightarrow{\text{treats}}$ cancers (Conf: 0.78, Rel: 0.70, Clr: 0.75)
@@ B.2 Knowledge graph from Proteomics articles (Generate using DeepSeek-V3, Examples)
B.3 Knowledge graph from Metabolomics articles (Generate using GLM-4, Examples)
Metabolomics Knowledge Graph Triples Key: Conf = Confidence, Rel = Relevance, Clr = Clarity G6PD $\xrightarrow{\text{activates}}$ NADPH (Conf: 0.75, Rel: 0.80, Clr: 0.50) BAG3 $\xrightarrow{\text{interacts\_with}}$ phosphorylation (Conf: 0.75, Rel: 0.50, Clr: 0.99) Mitotic NADPH surge $\xrightarrow{\text{prevents}}$ chromosome missegregation (Conf: 0.75, Rel: 0.80, Clr: 0.80) High BAG3 phosphorylation $\xrightarrow{\text{associated\_with}}$ poor prognosis (Conf: 0.75, Rel: 0.80, Clr: 0.99) G6PD $\xrightarrow{\text{crucial in}}$ pentose phosphate pathway (Conf: 0.85, Rel: 0.90, Clr: 0.80) Acetylation at lysine residue K89 $\xrightarrow{\text{activates}}$ G6PD (Conf: 0.75, Rel: 0.80, Clr: 0.99) Acetylation at lysine residue K403 $\xrightarrow{\text{inhibits}}$ G6PD (Conf: 0.75, Rel: 0.80, Clr: 0.80) astrocyte-to-neuron H2O2 signaling $\xrightarrow{\text{activates}}$ long-term memory formation (Conf: 0.75, Rel: 0.80, Clr: 0.80) astrocytes $\xrightarrow{\text{generates}}$ extracellular ROS (Conf: 0.75, Rel: 0.80, Clr: 0.80) extracellular ROS $\xrightarrow{\text{imported by}}$ neurons (Conf: 0.45, Rel: 0.80, Clr: 0.80) Alzheimer’s disease model $\xrightarrow{\text{impairs}}$ astrocyte-to-neuron H2O2 signaling (Conf: 0.75, Rel: 0.80, Clr: 0.80) Alzheimer’s disease model $\xrightarrow{\text{impairs}}$ memory formation (Conf: 0.85, Rel: 0.80, Clr: 0.80) ROS signaling $\xrightarrow{\text{important for}}$ memory (Conf: 0.75, Rel: 0.70, Clr: 0.60) astrocyte function $\xrightarrow{\text{important for}}$ memory (Conf: 0.75, Rel: 0.80, Clr: 0.60) Alzheimer’s disease $\xrightarrow{\text{involves}}$ astrocyte function (Conf: 0.75, Rel: 0.70, Clr: 0.50) Alzheimer’s disease $\xrightarrow{\text{involves}}$ ROS signaling (Conf: 0.75, Rel: 0.80, Clr: 0.80)

### v2

@@ 6 Conclusion
Acknowledge
This research was supported by National Key Research and Development Program of China (2024YFF0507400) and National Natural Science Foundation of China (6220071694).
Appendix
7 Limitations
@@
8 Detailed propmts for KARMA agents
@@
8.1 Function summaries of different agents
@@
8.2 Additional Notes on Prompt Engineering
@@
8.3 Ingestion Agent (IA) Prompt
@@
8.4 Reader Agent (RA) Prompt
@@
8.5 Summarizer Agent (SA) Prompt
@@
8.6 Entity Extraction Agent (EEA) Prompt
@@
8.7 Relationship Extraction Agent (REA) Prompt
@@
8.8 Schema Alignment Agent (SAA) Prompt
@@
8.9 Conflict Resolution Agent (CRA) Prompt
@@
8.10 Evaluator Agent (EA) Prompt for Confidence
@@
8.11 Evaluator Agent (EA) Prompt for Clarity
@@
8.12 Evaluator Agent (EA) Prompt for Relevance
@@
9 Examples of extracted knowledge graphs
9.1 Knowledge graph from Genomics articles (Generate using GPT-4o, Examples)
Genomics Knowledge Graph Triples Key: Conf = Confidence, Rel = Relevance, Clr = Clarity
EGFR $\xrightarrow{\text{causes}}$ lung adenocarcinoma (Conf: 0.75, Rel: 0.40, Clr: 0.60) EGFR $\xrightarrow{\text{causes}}$ non - small cell lung cancer (Conf: 0.85, Rel: 0.30, Clr: 0.70) T790M $\xrightarrow{\text{causes}}$ lung adenocarcinoma (Conf: 0.98, Rel: 0.20, Clr: 0.70) T790M $\xrightarrow{\text{causes}}$ non - small cell lung cancer (Conf: 0.95, Rel: 0.20, Clr: 0.80) EGFR $\xrightarrow{\text{activates}}$ EG (Conf: 0.75, Rel: 0.20, Clr: 0.99) EGFR $\xrightarrow{\text{causes}}$ proliferation (Conf: 0.85, Rel: 0.50, Clr: 0.60) MTX - 531 $\xrightarrow{\text{treats}}$ HNSCC (Conf: 0.99, Rel: 0.80, Clr: 0.50) MTX - 531 $\xrightarrow{\text{used\_in}}$ PDX (Conf: 0.75, Rel: 0.80, Clr: 0.50) PIK3CA mutations $\xrightarrow{\text{associated\_with}}$ HNSCC (Conf: 0.85, Rel: 0.70, Clr: 0.99) MTX - 531 $\xrightarrow{\text{treats}}$ PIK3CA mutations (Conf: 0.65, Rel: 0.40, Clr: 0.99) MCLA - 158 $\xrightarrow{\text{targets}}$ EGFR (Conf: 0.75, Rel: 0.30, Clr: 0.99) EGFR $\xrightarrow{\text{interacts\_with}}$ MCLA - 158 (Conf: 0.75, Rel: 0.20, Clr: 0.80) PDOs $\xrightarrow{\text{used\_in}}$ biobank (Conf: 0.75, Rel: 0.50, Clr: 0.50) Retinoic acid $\xrightarrow{\text{disrupts}}$ autocrine growth pathway (Conf: 0.75, Rel: 0.80, Clr: 0.60) Retinoic acid $\xrightarrow{\text{interacts\_with}}$ nuclear retinoic acid receptors (Conf: 0.95, Rel: 0.80, Clr: 0.99) nuclear retinoic acid receptors $\xrightarrow{\text{regulate}}$ gene transcription (Conf: 0.95, Rel: 0.70, Clr: 0.50) EGFR $\xrightarrow{\text{interacts\_with}}$ Cys797 (Conf: 0.95, Rel: 0.80, Clr: 0.99)
@@ 9.1 Knowledge graph from Genomics articles (Generate using GPT-4o, Examples)
9.2 Knowledge graph from Proteomics articles (Generate using DeepSeek-V3, Examples)
Proteomics Knowledge Graph Triples Key: Conf = Confidence, Rel = Relevance, Clr = Clarity
p53 $\xrightarrow{\text{induces}}$ cell-cycle arrest (Conf: 0.95, Rel: 0.70, Clr: 0.85) p53 $\xrightarrow{\text{induces}}$ apoptosis (Conf: 0.95, Rel: 0.70, Clr: 0.85) mutant p53 $\xrightarrow{\text{causes}}$ chemotherapy resistance (Conf: 0.85, Rel: 0.85, Clr: 0.85) MDM2 $\xrightarrow{\text{interacts\_with}}$ p53 (Conf: 0.95, Rel: 0.20, Clr: 0.90) PRIMA-1 $\xrightarrow{\text{induces}}$ apoptosis (Conf: 0.85, Rel: 0.70, Clr: 0.85) NOS2 $\xrightarrow{\text{associated\_with}}$ cancers (Conf: 0.85, Rel: 0.70, Clr: 0.60) NOS2 inhibitors $\xrightarrow{\text{inhibits}}$ NOS2 (Conf: 0.95, Rel: 0.80, Clr: 0.90) NOS2 inhibitors $\xrightarrow{\text{reduces}}$ tumor growth (Conf: 0.85, Rel: 0.75, Clr: 0.85) NO $\xrightarrow{\text{induces}}$ VEGF (Conf: 0.78, Rel: 0.70, Clr: 0.70) NO $\xrightarrow{\text{induces}}$ neovascularization (Conf: 0.45, Rel: 0.70, Clr: 0.75) NOS2 inhibitors $\xrightarrow{\text{has\_therapeutic\_potential}}$ p53-mutant cancers (Conf: 0.78, Rel: 0.75, Clr: 0.85) tumor progression $\xrightarrow{\text{dependent\_on}}$ p53 (Conf: 0.85, Rel: 0.70, Clr: 0.85) NOS2 $\xrightarrow{\text{promotes}}$ tumor growth (Conf: 0.85, Rel: 0.85, Clr: 0.75) NOS2 $\xrightarrow{\text{produces}}$ nitric oxide (NO) (Conf: 0.95, Rel: 0.90, Clr: 0.90) hypoxia $\xrightarrow{\text{regulates}}$ iNOS expression (Conf: 0.85, Rel: 0.70, Clr: 0.85) iNOS expression $\xrightarrow{\text{influences}}$ endothelial integrity (Conf: 0.85, Rel: 0.70, Clr: 0.75) sulindac sulfide $\xrightarrow{\text{treats}}$ cancers (Conf: 0.78, Rel: 0.70, Clr: 0.75)
@@ 9.2 Knowledge graph from Proteomics articles (Generate using DeepSeek-V3, Examples)
9.3 Knowledge graph from Metabolomics articles (Generate using GLM-4, Examples)
Metabolomics Knowledge Graph Triples Key: Conf = Confidence, Rel = Relevance, Clr = Clarity
G6PD $\xrightarrow{\text{activates}}$ NADPH (Conf: 0.75, Rel: 0.80, Clr: 0.50) BAG3 $\xrightarrow{\text{interacts\_with}}$ phosphorylation (Conf: 0.75, Rel: 0.50, Clr: 0.99) Mitotic NADPH surge $\xrightarrow{\text{prevents}}$ chromosome missegregation (Conf: 0.75, Rel: 0.80, Clr: 0.80) High BAG3 phosphorylation $\xrightarrow{\text{associated\_with}}$ poor prognosis (Conf: 0.75, Rel: 0.80, Clr: 0.99) G6PD $\xrightarrow{\text{crucial in}}$ pentose phosphate pathway (Conf: 0.85, Rel: 0.90, Clr: 0.80) Acetylation at lysine residue K89 $\xrightarrow{\text{activates}}$ G6PD (Conf: 0.75, Rel: 0.80, Clr: 0.99) Acetylation at lysine residue K403 $\xrightarrow{\text{inhibits}}$ G6PD (Conf: 0.75, Rel: 0.80, Clr: 0.80) astrocyte-to-neuron H2O2 signaling $\xrightarrow{\text{activates}}$ long-term memory formation (Conf: 0.75, Rel: 0.80, Clr: 0.80) astrocytes $\xrightarrow{\text{generates}}$ extracellular ROS (Conf: 0.75, Rel: 0.80, Clr: 0.80) extracellular ROS $\xrightarrow{\text{imported by}}$ neurons (Conf: 0.45, Rel: 0.80, Clr: 0.80) Alzheimer’s disease model $\xrightarrow{\text{impairs}}$ astrocyte-to-neuron H2O2 signaling (Conf: 0.75, Rel: 0.80, Clr: 0.80) Alzheimer’s disease model $\xrightarrow{\text{impairs}}$ memory formation (Conf: 0.85, Rel: 0.80, Clr: 0.80) ROS signaling $\xrightarrow{\text{important for}}$ memory (Conf: 0.75, Rel: 0.70, Clr: 0.60) astrocyte function $\xrightarrow{\text{important for}}$ memory (Conf: 0.75, Rel: 0.80, Clr: 0.60) Alzheimer’s disease $\xrightarrow{\text{involves}}$ astrocyte function (Conf: 0.75, Rel: 0.70, Clr: 0.50) Alzheimer’s disease $\xrightarrow{\text{involves}}$ ROS signaling (Conf: 0.75, Rel: 0.80, Clr: 0.80)